# Without Paired Labeled Data: End-to-End Self-Supervised Learning for Drone-view Geo-Localization

### v3

@@ Abstract
Drone-view Geo-Localization (DVGL) aims to achieve accurate localization of drones by retrieving the most relevant GPS-tagged satellite images. However, most existing methods heavily rely on strictly pre-paired drone-satellite images for supervised learning. When the target region shifts, new paired samples are typically required to adapt to the distribution changes. The high cost of annotation and the limited transferability of these methods significantly hinder the practical deployment of DVGL in open-world scenarios. To address these limitations, we propose an end-to-end self-supervised learning method with a shallow backbone network. It employs a clustering algorithm to generate pseudo-labels and adopts a dual-path contrastive learning framework to learn discriminative intra-view representations. Furthermore, our method incorporates two core modules, including the dynamic hierarchical memory learning module (DHML) and the information consistency evolution learning module (ICEL). The DHML combines short-term and long-term memory to enhance intra-view feature consistency and discriminability. Meanwhile, the ICEL module utilizes a neighborhood-driven dynamic constraint mechanism to systematically capture implicit cross-view semantic correlations, consequently improving cross-view feature alignment. To further stabilize and strengthen the self-supervised training process, a pseudo-label enhancement strategy is introduced to enhance the quality of pseudo supervision. Extensive experiments on three public benchmark datasets demonstrate that the proposed method consistently outperforms existing self-supervised methods and even surpasses several state-of-the-art supervised methods. Our code is available at https://github.com/ISChenawei/DMNIL.
Index Terms:
Drone-view Geo-Localization, Self-Supervised Learning, Contrastive Learning.
@@ I Introduction
Drone-View Geo-Localization (DVGL) is intended to achieve high-precision localization of drones in GPS-denied scenarios by retrieving GPS-tagged satellite images that correspond to drone aerial images [1]. This capability is critical for applications such as autonomous navigation, disaster rescue, and emergency response. In recent years, existing DVGL methods have mainly relied on supervised learning using strictly paired drone-satellite images [2, 3]. Despite the promising performance demonstrated by these supervised methods, several intrinsic limitations are exposed. Firstly, the acquisition of instance-level paired drone-satellite images incurs additional training expenses [4]. Secondly, large amounts of unpaired cross-view image resources remain under-exploited. Lastly, when these methods are applied to new scenarios, the need arises for re-collection and re-pairing of drone-satellite images for training purposes to accommodate domain shifts. These limitations significantly hinder the practical deployment of DVGL in open-world scenarios. Thereby, it is necessary to explore a self-supervised method that can extract latent cross-view correlations from large-scale unpaired drone-satellite images.
Figure 1: As representative cross-view retrieval tasks, ReID and DVGL share the common objective of establishing correspondences between query and gallery images based on discriminative feature representations. The notable progress achieved by ReID, especially in self-supervised learning scenarios, offers a valuable methodological foundation and insightful paradigms for advancing DVGL research. However, DVGL presents greater challenges for effective feature learning and discrimination compared with ReID. (A) ReID tasks are characterized by relatively small intra-class variations and distinct inter-class differences, which facilitates effective feature discrimination. (B) In contrast, DVGL tasks suffer from large intra-class variations and small inter-class differences, leading to significant cross-view gaps and intra-view ambiguities that hinder feature discrimination.
@@
In recent years, self-supervised person re-identification (ReID) methods [5, 6] have achieved remarkable progress to offer crucial theoretical foundations for DVGL. Typically, these methods employ clustering algorithms [7, 8] on features extracted by a backbone network to generate pseudo-labels, which serve as the initial supervision signals for self-supervised learning. However, as illustrated in Fig. 1, compared to conventional ReID tasks, DVGL presents more severe challenges, such as large intra-view ambiguities and cross-view discrepancies. As shown in Fig. 2, these issues hinder the backbone network from extracting stable and discriminative features at the early stage, which directly undermines the quality of pseudo-labels generated during the initial clustering phase. Without effective strategies to guide the learning of intra-view and cross-view consistency as well as feature discrimination, the model lacks the necessary ability to extract meaningful representations from low-quality pseudo-labels. As a result, the learning process may precipitate a detrimental feedback loop, wherein inaccurate pseudo-labels undermine the quality of feature representations to further compromise pseudo-label quality in subsequent iterations, ultimately leading to the collapse of self-supervised training. Although a current self-supervised method [9] proposed a coarse-to-fine framework for DVGL with a similar dual-path contrastive learning framework, it relies on explicit appearance-level transformations rather than intrinsically modeling the underlying cross-view feature relationships. In the shared space, it still exhibits limited cross-view alignment and susceptibility to feature noise.
Figure 2: Comparison of Recall@k performance on the DVGL and ReID datasets using ConvNeXt-Tiny [10], which is pre-trained on ImageNet-22k [11]. The evaluation is conducted on the DVGL dataset University-1652 [12] and the four person ReID datasets including MSMT17 [13], DukeMTMC-reID [14], CUHK03 [15], and Market1501 [16].
@@
In this paper, we propose a novel end-to-end self-supervised learning method for DVGL. A lightweight and shallow ConvNeXt-Tiny backbone network [10] is employed to extract features, which are subsequently clustered to generate pseudo-labels. Combined with a dual-path contrastive learning strategy [17, 18], the backbone network coupled with a downstream clustering process forms the baseline network of our method for learning intra-view feature representations. Building upon this baseline, we design the dynamic hierarchical memory learning (DHML) module and the information consistency evolution learning (ICEL) module as two key components of our method. The DHML module captures the relationship between short-term and long-term memory to enhance the discriminability of intra-view feature representations. The ICEL module constructs a neighborhood-driven dynamic constraint mechanism by integrating neighborhood consistency constraints with mutual information optimization to facilitate cross-view feature alignment and discriminative representation learning. DHML enhances intra-view discrimination and ICEL improves cross-view consistency. Through joint optimization of intra-view and cross-view features, they collectively strengthen the feature discriminability and consistency. Additionally, a pseudo-label enhancement (PLE) strategy is employed to improve the reliability of pseudo-labels and guide more stable self-supervised feature learning. The main contributions can be summarized as follows.
@@
  • We propose an end-to-end self-supervised learning method for DVGL. Built upon a shallow backbone and a dual-path contrastive learning strategy, our method integrates dynamic hierarchical memory learning and information consistency evolution learning to enhance feature consistency and discriminability for both intra-view and cross-view representation learning. Additionally, a pseudo-label enhancement (PLE) strategy is employed to guide more stable self-supervised learning.
@@
  • Extensive experiments demonstrate that our method, which employs a lightweight backbone, outperforms existing self-supervised methods on three DVGL benchmarks and even surpasses several state-of-the-art supervised methods.
@@ II Related Works
This section establishes a systematic and comprehensive theoretical framework to support the development of the proposed self-supervised learning method for DVGL. DVGL is typically regarded as a sub-task of cross-view geo-localization (CVGL). To provide a more comprehensive understanding of the field, we first review the overall progress in CVGL, including both supervised and self-supervised methods. Then, we summarize the main methods in self-supervised ReID that have provided analogous problem-solving paradigms.
@@ II-A Supervised Cross-View Geo-localization
CVGL aims to achieve high-precision localization of query images by retrieving GPS-tagged satellite images. Early research focused mainly on the CVGL in ground-view [19] with several benchmark datasets, such as CVUSA [20], CVACT [21], and VIGOR [22]. To address the challenge of ground-view CVGL, researchers have proposed various methods. CVM-Net [23] employed NetVLAD [24] to encode images as global descriptors to reduce visual discrepancies caused by viewpoint variations. TransGeo [25] utilized attention-based zooming and sharpness-aware optimization to enhance detail learning and improve model generalization. GeoDTR [26] leveraged a Transformer-based [27] feature extractor to disentangle geometric features, effectively addressing perspective shifts. Sample4Geo [28] introduced a hard negative mining strategy, focusing on challenging negative samples to further improve discriminative capability.
With the continuous development and application of drone technology, DVGL has gradually become an important research direction and has achieved significant progress [29, 30, 31]. Several benchmark datasets for DVGL have been proposed, such as University-1652 [12], SUES-200 [32], and DenseUAV [33], which have driven further research in this field. In this scenario, researchers have proposed a variety of methods. LPN [34] extracted fine-grained features through local pattern partitioning to improve the ability to capture local details. IFS [35] utilized a multi-branch strategy to effectively fuse global and local features. CAMP [29] improved the accuracy of cross-view matching by contrastive attribute mining and position-aware partitioning, while DAC [3] enhanced feature consistency and matching accuracy by introducing domain alignment and scene consistency constraints. MEAN [2] adopted progressive embedding diversification, global-to-local associations, and cross-domain enhanced alignment to further improve the representation and alignment of features.
Nevertheless, the remarkable performances exhibited by these DVGL methods require extensive pre-processed drone-satellite image pairs. In this work, we concentrate on self-supervised learning approaches for DVGL, particularly addressing scenarios where labeled pairings are unavailable, which is crucial for real-world applications.
@@ II-B Self-Supervised Cross-View Geo-localization
Due to the dependence of CVGL on extensive pre-processed image pairs, which significantly increases costs and limits applications for the real world, researchers have begun to explore self-supervised learning methods for CVGL. The work in [36] proposed a self-supervised learning method that used correspondence-free projection to transform ground panorama images into bird’s eye view images. It also applied CycleGAN [37] to generate fake satellite images, coupled with an additional SAFA module [38] to reduce discrepancies. However, this explicit alignment method has inherent limitations in capturing the deep semantic relationships of cross-view features and may introduce noise. Furthermore, the non-end-to-end training method employed significantly increases training time and computational costs with low practical applicability. Additionally, existing work [4] utilized frozen foundation models (FMs) [39] to extract features, and a self-supervised adapter was introduced to mitigate visual discrepancies between different points of view. However, this method heavily relied on the ability of FMs for feature extraction. When FMs fail to effectively handle viewpoint discrepancies, the performance of the adapter is also affected. To reduce reliance on the feature extractor, CDIKTNet [40] trained with a small amount of paired drone-satellite images to optimize the initial feature distribution and then performed clustering to generate high-quality pseudo-labels. However, this method inherently depends on the feature extractor, as it still requires limited paired drone-satellite images for effective learning.
The above methods have made meaningful exploratory attempts. However, these methods still face challenges, such as the additional computation time and overhead introduced by non-end-to-end frameworks, as well as dependence on the capability of feature extractor. In contrast, our method is fully end-to-end, eliminating the need for additional training processes and completely removing the reliance on feature extractor initialization.
@@ II-C Methodological Parallels from Self-Supervised ReID
ReID [41] aims to match the same person image captured by non-overlapping cameras. Due to its critical role in video surveillance, ReID has attracted extensive research attention and achieved remarkable progress. It has provided a fully end-to-end framework to eliminate the need for additional training processes and completely remove the reliance on feature extractor initialization, which offers strong inspiration for our consequent work in the DVGL task. However, these achievements are facilitated by extensive human-labeled data. To address this limitation, self-supervised learning methods [5, 6] have been proposed. Among them, clustering algorithms are widely employed to generate pseudo-labels, which serve as supervision signals for model training. In addition, benefitting from the introduction of instance-level contrastive learning [42] with memory banks and momentum update strategies [43], many methods [44, 45] consider each unlabeled sample as an individual class to learn discriminative instance-level feature representations. However, these instance-level features were largely modified during training, which resulted in unstable representations [46]. To address this issue, recent studies [47, 48] have begun exploring cluster-level relationships to discover more stable and robust associations among pedestrian representations. Cluster-Contrast [17] contrasts instance features with these cluster-level representations to address the issue of inconsistent instance discrimination. Nevertheless, in cross-modal ReID scenarios, where images are captured by heterogeneous sensors, modality discrepancies hinder the effectiveness of these methods in learning modality-invariant features. ADCA [18] introduced a dual-path contrastive framework to bridge modality gaps and enhance cross-modal feature alignment. Building on this, SDCL [49] employed a dual-path Transformer for shallow-to-deep contrastive learning, and refined pseudo-labels via consistency constraints across feature hierarchies to achieve more stable and discriminative cross-modal representations. The above methods focus on alleviating the dependency on human-labeled data in the ReID and provide valuable insights for DVGL. However, DVGL suffers from more severe viewpoint and scale variations. Furthermore, the spatial continuity of geographic environments inevitably induces feature ambiguity, which would severely degrade clustering effectiveness.
@@
Figure 3: The overall pipeline of our method consists of a lightweight backbone, a dual-path contrastive learning strategy, a dynamic hierarchical memory learning module, and an information consistency evolution module. Specifically, the dual-path contrastive learning is designed to learn discriminative and consistent intra-view feature representations. The dynamic hierarchical memory module further captures intra-view feature variations under different perspectives and scales, thereby enhancing the robustness and discriminability of the learned representations. The information consistency evolution module focuses on modeling cross-view feature consistency through a neighborhood-driven learning strategy, and further improves the training process by integrating a pseudo-label enhancement strategy.

### v5

@@ Abstract
Drone-view Geo-Localization (DVGL) aims to achieve accurate localization of drones by retrieving the most relevant GPS-tagged satellite images. However, most existing methods heavily rely on strictly pre-paired drone-satellite images for supervised learning. When the target region shifts, new paired samples are typically required to adapt to the distribution changes. The high cost of annotation and the limited transferability of these methods significantly hinder the practical deployment of DVGL in open-world scenarios. To address these limitations, we propose a novel end-to-end self-supervised learning method with a shallow backbone network, called the dynamic memory-driven and neighborhood information learning (DMNIL) method. It employs a clustering algorithm to generate pseudo-labels and adopts a dual-path contrastive learning framework to learn discriminative intra-view representations. Furthermore, DMNIL incorporates two core modules, including the dynamic hierarchical memory learning (DHML) module and the information consistency evolution learning (ICEL) module. The DHML module combines short-term and long-term memory to enhance intra-view feature consistency and discriminability. Meanwhile, the ICEL module utilizes a neighborhood-driven dynamic constraint mechanism to systematically capture implicit cross-view semantic correlations, consequently improving cross-view feature alignment. To further stabilize and strengthen the self-supervised training process, a pseudo-label enhancement strategy is introduced to enhance the quality of pseudo supervision. Extensive experiments on three public benchmark datasets demonstrate that the proposed method consistently outperforms existing self-supervised methods and even surpasses several state-of-the-art supervised methods. Our code is available at https://github.com/ISChenawei/DMNIL.
@@ I Introduction
Drone-View Geo-Localization (DVGL) is intended to achieve high-precision localization of drones in GPS-denied scenarios by retrieving GPS-tagged satellite images that correspond to drone aerial images [shen2023mccg]. This capability is critical for applications such as autonomous navigation, disaster rescue, and emergency response. In recent years, existing DVGL methods have mainly relied on supervised learning using strictly paired drone-satellite images [chen2024multi, xia2024enhancing]. Despite the promising performance demonstrated by these supervised methods, several intrinsic limitations are exposed. Firstly, the acquisition of instance-level paired drone-satellite images incurs additional training expenses [li2024learning]. Despite the promising performance demonstrated by these supervised methods, several intrinsic limitations are exposed. Firstly, the acquisition of instance-level paired drone-satellite images incurs additional training expenses. Automatically pairing drone-satellite images using GPS is infeasible due to unreliable GPS signals and severe viewpoint discrepancies between drone and satellite imagery, so existing datasets [zheng2020university, zhu2023sues, wang2024multiple] rely heavily on manual screening and multi-stage human verification to ensure accurate cross-view. Secondly, large amounts of unpaired cross-view image resources remain under-exploited. Lastly, when these methods are applied to new scenarios, the need arises for re-collection and re-pairing of drone-satellite images for training purposes to accommodate domain shifts. These limitations hinder the practical deployment of DVGL. Although researchers have begun to explore self-supervised DVGL methods to address these challenges, existing methods remain limited. The existing method [chen2025limited] attempted to discover cross-view relationships from unpaired data through multi-stage training initialized with a small amount of paired data, as shown in Fig. 1(A(a)). As shown in Fig. 1(A(b)), concurrent work [wang2025coarse] attempted to use intermediate states to bridge the cross-view gap. In contrast, our method aims to directly and effectively uncover latent cross-view correlations in an end-to-end manner without intermediate states, as illustrated in Fig. 1(B).
@@
In recent years, self-supervised person re-identification (ReID) methods [yin2023real, 10680455] have achieved remarkable progress to offer crucial theoretical foundations for DVGL. Typically, these methods employ clustering algorithms [ester1996density, macqueen1967some] on features extracted by a backbone network to generate pseudo-labels, which serve as the initial supervision signals for self-supervised learning. The clustering-based pseudo-label generation solution has been widely adopted and has proven highly effective in self-supervised ReID. Inspired by this, we adopt a similar solution in our method to obtain pseudo-labels. However, as illustrated in Fig. 1(C) and Fig. 1(D), compared with conventional ReID tasks, DVGL presents more severe challenges, such as large intra-view ambiguities and cross-view discrepancies. As shown in Fig. 1(E), these issues hinder the backbone network from extracting stable and discriminative features at the early stage, which directly undermines the quality of pseudo-labels generated during the initial clustering phase. To address these, it would be essential to design effective strategies to guide drone features and satellite features to achieve effective intra-class cohesion and inter-class separation within their respective views, which can form distinct clusters. Simultaneously, cross-view consistency learning process would be deliberately optimized to preserve cross-view feature correlations. In the absence of such guidance, the model lacks the necessary ability to learn meaningful representations from low-quality pseudo-labels. This deficiency can trigger a degenerative feedback loop, wherein inaccurate pseudo-labels undermine the quality of feature representations to further compromise pseudo-label quality in subsequent iterations, ultimately leading to the collapse of self-supervised training.
Figure 1: Motivation, Challenges, and Evaluation. (A-B): Motivation of the proposed DMNIL, which aims to explore the latent relationships between cross-view features, without relying on any paired data and multi-stage training and seeking intermediate transitional states to bridge the cross-view gap. (C-D): As representative cross-view retrieval tasks, ReID and DVGL share the common objective of establishing correspondences between query and gallery images based on discriminative feature representations. The notable progress achieved by ReID, especially in self-supervised learning scenarios, offers a valuable methodological foundation and insightful paradigms for advancing DVGL research. However, DVGL presents greater challenges for effective feature learning and discrimination compared to ReID. (C): ReID tasks are characterized by relatively small intra-class variations and distinct inter-class differences, which facilitates effective feature discrimination. (D): In contrast, DVGL tasks suffer from large intra-class variations and small inter-class differences, leading to cross-view gaps and intra-view ambiguities that hinder feature discrimination. (E): The Recall@k performance of the solely used backbone ConvNeXt-Tiny [liu2022convnet] that is pre-trained on ImageNet-22k[deng2009imagenet] is presented on two benchmarks, where the evaluation is conducted on the DVGL dataset University-1652[zheng2020university] and the four person ReID datasets including MSMT17[wei2018person], DukeMTMC-reID[zheng2017discriminatively], CUHK03[sun2014deep], and Market1501[zheng2015scalable]. (F): Performance of DMNIL is compared with state-of-the-art self-supervised and supervised methods, which outperforms existing state-of-the-art self-supervised methods and even surpasses several supervised methods.
@@
In this paper, we propose a novel end-to-end self-supervised learning method for DVGL, namely the dynamic memory-driven and neighborhood information learning (DMNIL) method. A lightweight and shallow ConvNeXt-Tiny backbone network [liu2022convnet] is employed to extract features, which are subsequently clustered to generate pseudo-labels. Combined with a dual-path contrastive learning strategy [dai2022cluster, adca], the backbone network coupled with a downstream clustering process forms the baseline network of DMNIL for learning intra-view feature representations. Building upon this baseline, we design the dynamic hierarchical memory learning (DHML) module and the information consistency evolution learning (ICEL) module as two key components of DMNIL. The DHML module captures the relationship between short-term and long-term memory to enhance the discriminability of intra-view feature representations. The ICEL module constructs a neighborhood-driven dynamic constraint mechanism by integrating neighborhood consistency constraints with mutual information optimization to facilitate cross-view feature alignment and discriminative representation learning. DHML enhances intra-view discrimination and ICEL improves cross-view consistency. Through joint optimization of intra-view and cross-view features, they collectively strengthen the feature discriminability and consistency. Additionally, a pseudo-label enhancement (PLE) strategy is employed to improve the reliability of pseudo-labels and guide more stable self-supervised feature learning. DMNIL surpasses the state-of-the-art self-supervised approaches and achieves performance comparable to supervised methods, as illustrated in Fig. 1(F). The main contributions can be summarized as follows.
@@
  • We propose an effective dynamic memory-driven and neighborhood information learning method for self-supervised DVGL. Built upon a shallow backbone and a dual-path contrastive learning strategy, our method integrates dynamic hierarchical memory learning and information consistency evolution learning to enhance feature consistency and discriminability for both intra-view and cross-view representation learning. Additionally, a PLE strategy is employed to guide more stable self-supervised learning.
@@
  • Extensive experiments demonstrate that DMNIL, which employs a lightweight backbone, outperforms existing self-supervised methods on three benchmarks and even surpasses several state-of-the-art supervised methods.
@@ II Related Works
This section establishes a systematic and comprehensive theoretical framework to support the development of the self-supervised learning method proposed for DVGL. DVGL is typically regarded as a sub-task of cross-view geo-localization (CVGL). To provide a more comprehensive understanding of the field, we first review the overall progress in CVGL, including both supervised and self-supervised methods. Then, we summarize the main methods in self-supervised ReID that have provided analogous problem-solving paradigms.
@@ II-A Supervised Cross-View Geo-localization
CVGL aims to achieve high-precision localization of query images by retrieving GPS-tagged satellite images. Early research focused mainly on the CVGL in ground-view [10601183] with several benchmark datasets, such as CVUSA [zhai2017predicting], CVACT [2019Lending], and VIGOR [zhu2021vigor]. To address the challenge of ground-view CVGL, researchers have proposed various methods. CVM-Net [Hu_2018_CVPR] employed NetVLAD [arandjelovic2016netvlad] to encode images as global descriptors to reduce visual discrepancies caused by viewpoint variations. TransGeo [zhu2022transgeo] utilized attention-based zooming and sharpness-aware optimization to enhance detail learning and improve model generalization. GeoDTR [zhang2023cross] leveraged a Transformer-based [liu2022swin] feature extractor to disentangle geometric features, which effectively address cross-view shifts. Sample4Geo [deuser2023sample4geo] introduced a hard negative mining strategy, focusing on challenging negative samples to further improve discriminative capability.
With the continuous development and application of drone technology, DVGL has gradually become an important research direction and has achieved significant progress [wu2024camp, lv2024direction, du2024ccr]. Several benchmark datasets for DVGL have been proposed, such as University-1652 [zheng2020university], SUES-200 [zhu2023sues], and DenseUAV [DenseUAV], which have driven further research in this field. In this scenario, researchers have proposed a variety of methods. LPN [wang2021each] extracted fine-grained features through local pattern partitioning to improve the ability to capture local details. IFS [ge2024multibranch] utilized a multi-branch strategy to effectively fuse global and local features. CAMP [wu2024camp] improved the accuracy of cross-view matching by contrastive attribute mining and position-aware partitioning, while DAC [xia2024enhancing] enhanced feature consistency and matching accuracy by introducing domain alignment and scene consistency constraints. MEAN [chen2024multi] adopted progressive embedding diversification, global-to-local associations, and cross-domain enhanced alignment to further improve the representation and alignment of features.
Nevertheless, the remarkable performances exhibited by these DVGL methods require extensive pre-processed drone-satellite image pairs. In this work, we concentrate on self-supervised learning method for DVGL, particularly addressing scenarios where labeled pairings are unavailable, which is crucial for real-world applications.
@@ II-B Self-Supervised Cross-View Geo-localization
Due to the dependence of CVGL on extensive pre-processed image pairs, which significantly increases costs and limits applications for the real world, researchers have begun to explore self-supervised learning methods for CVGL. The work in [Li_2024_CVPR] proposed a self-supervised learning method that used correspondence-free projection to transform ground panorama images into bird’s eye view images. It also applied CycleGAN [Zhu_2017_ICCV] to generate fake satellite images, coupled with an additional SAFA module [shi2019spatial] to reduce discrepancies. However, this explicit alignment method has inherent limitations in capturing the deep semantic relationships of cross-view features and may introduce noise. Furthermore, the non-end-to-end training method employed increases training time and computational costs with low practical applicability. Additionally, existing work [li2024learning] utilized frozen foundation models (FMs) [oquab2023dinov2] to extract features, and a self-supervised adapter was introduced to mitigate visual discrepancies between different points of view. However, this method heavily relied on the ability of FMs for feature extraction. When FMs fail to effectively handle viewpoint discrepancies, the performance of the adapter is also affected. To reduce reliance on the feature extractor, CDIKTNet [chen2025limited] trained with a small amount of paired drone-satellite images to optimize the initial feature distribution and then performed clustering to generate high-quality pseudo-labels. However, this method inherently depends on the feature extractor, as it still requires limited paired drone-satellite images for effective learning.
The above methods have made meaningful exploratory attempts. However, these methods still face challenges, such as the additional computation time and overhead introduced by non-end-to-end frameworks, as well as dependence on the capability of feature extractor. In contrast, DMNIL is fully end-to-end, eliminating the need for additional training processes and completely removing the reliance on feature extractor initialization.
Figure 2: Pipeline Overview. The DMNIL consists of a lightweight backbone, a dual-path contrastive learning strategy, a dynamic hierarchical memory learning module, and an information consistency evolution module. Specifically, the dual-path contrastive learning is designed to learn discriminative and consistent intra-view feature representations. The dynamic hierarchical memory module further captures intra-view feature variations under different viewpoints and scales, thereby enhancing the robustness and discriminability of the learned representations. The information consistency evolution module focuses on modeling cross-view feature consistency through a neighborhood-driven learning strategy, and further improves the training process by integrating a pseudo-label enhancement strategy.
@@ II-C Methodological Parallels from Self-Supervised ReID
ReID [9457243, 10310270] aims to match the same person image captured by non-overlapping cameras. Due to its critical role in video surveillance, ReID has attracted extensive research attention and achieved remarkable progress. It has provided a fully end-to-end framework to eliminate the need for additional training processes and completely remove the reliance on feature extractor initialization, which offers strong inspiration for our consequent work in the DVGL task. However, these achievements are facilitated by extensive human-labeled data. To address this limitation, self-supervised learning methods [yin2023real, 9840394] have been proposed. Among them, clustering algorithms are widely employed to generate pseudo-labels, which serve as supervision signals for model training. In addition, benefitting from the introduction of instance-level contrastive learning [Wu_2018_CVPR] with memory banks and momentum update strategies [He_2020_CVPR], many methods [chen2021ice, wang2022optimal] consider each unlabeled sample as an individual class to learn discriminative instance-level feature representations. However, these instance-level features were largely modified during training, which resulted in unstable representations [10882953]. To address this issue, recent studies [Yang_2023_ICCV, cheng2023efficient] have begun exploring cluster-level relationships to discover more stable and robust associations among pedestrian representations. Cluster-Contrast [dai2022cluster] contrasts instance features with these cluster-level representations to address the issue of inconsistent instance discrimination. Nevertheless, in cross-modal ReID scenarios, where images are captured by heterogeneous sensors, modality discrepancies hinder the effectiveness of these methods in learning modality-invariant features. ADCA [adca] introduced a dual-path contrastive framework to bridge modality gaps and enhance cross-modal feature alignment. Building on this, SDCL [yang2024shallow] employed a dual-path Transformer for shallow-to-deep contrastive learning, and refined pseudo-labels via consistency constraints across feature hierarchies to achieve more stable and discriminative cross-modal representations. The above methods focus on alleviating the dependency on human-labeled data in the ReID and provide valuable insights for DVGL. However, DVGL suffers from more severe viewpoint and scale variations. Furthermore, the spatial continuity of geographic environments inevitably induces feature ambiguity, which would severely degrade clustering effectiveness.